I am having trouble understanding how to solve systems of inequalities. Here is an example:
8x-7y+-3
6x-5y=-1
How do I solve each system (using any method)
Thanks,
Carol

Answer by checkley79(3341)

8x-7y=-3 MULTIPLY BY 3. [NOTE: I CHANGED THE + TO AN = SIGN]
6x-5y=-1 MULTIPLY BY -4 & ADD.
24X-21Y=-9
-24X+20Y=4 ADD.
---------------------
-Y=-5
Y=5 ANS.
8X-7*5=-3
8X-35=-3
8X=-3+35
8X=32
X=32/8
X=4 ANS.
PROOF:
6*4-5*5=-1
24-25=-1
-1=-1